User name markowitz
Log entry time 23:57:11 on February 4,2002
Entry number 81805
completed change to kin 5c. Set timing to 2.890
2062 9-foil optics 100K events
16:39 end run 2062 at 66k - mcc called
to say beam is off for 1/2 hr to do beam tuning
2066,2067 LH2, magnet off, no radiator
turning on magnet, and adjusting threshold to 1 GeV
2068 LH2, magnet on, no radiator
putting in the radiator and changing the angle to 32.66. Adjusting the momentum 2.772 GeV.
2069 LH2, magnet on, radiator in
timing peak did not show, so we set to 2.772
2070 timing maybe there after all
2074 same but end-of-run failed. CODA needed rebooted twice.
2077 same but did not seem to get the correct run number. We reset the run number to 2080 (so 2078 and 2079 and 2080 do not exist on tape).
Estimate is that we need .4 C to get the statistics here at 5C.